### v1

# The Influence of Human-inspired Agentic Sophistication in LLM-driven Strategic Reasoners

Abstract
The rapid rise of large language models (LLMs) has shifted artificial intelligence (AI) research toward agentic systems, motivating the use of weaker and more flexible notions of agency. However, this shift raises key questions about the extent to which LLM-based agents replicate human strategic reasoning, particularly in game-theoretic settings. In this context, we examine the role of agentic sophistication in shaping artificial reasoners’ performance by evaluating three agent designs: a simple game-theoretic model, an unstructured LLM-as-agent model, and an LLM integrated into a traditional agentic framework. Using guessing games as a testbed, we benchmarked these agents against human participants across general reasoning patterns and individual role-based objectives. Furthermore, we introduced obfuscated game scenarios to assess agents’ ability to generalise beyond training distributions. Our analysis, covering over 2000 reasoning samples across 25 agent configurations, shows that human-inspired cognitive structures can enhance LLM agents’ alignment with human strategic behaviour. Still, the relationship between agentic design complexity and human-likeness is non-linear, highlighting a critical dependence on underlying LLM capabilities and suggesting limits to simple architectural augmentation.
[…]
7846
1 Introduction
The transition from LLMs to agentic AI prompts a fundamental reconsideration of traditional models of agency and multi-agent systems [18]. Earlier conceptions of agency [46] typically invoked the intentional stance, framing agents as belief holding entities that are capable of deliberation and reasoning with them to generate and enact their intentions [4, 32]. In contrast, the growing success of LLMs has led to their increasing characterisation as inherently agentic [23] – evaluated not only as capable of reasoning, but also as demonstrating cognitive abilities [1, 37, 40], such as forming recursive (or nested) beliefs about others’ mental states, as if possessing a theory of mind [45], much like humans [35, 9].
As part of a broader research agenda on using agentic AI for game-theoretic social simulations [36], we are interested in developing agentic AIs exhibiting human-like strategic reasoning for navigating scenarios of varying complexity, anticipating others’ actions, and selecting responses accordingly [33]. Experimental game theory has yielded extensive evidence documenting human performance in such recursive reasoning tasks [8, 11, 12, 14, 28, 29]. Building on this foundation, studies have begun to examine the capabilities of LLMs in competitive environments [27, 26, 24, 10, 48, 13]. Although agentic applications are proliferating rapidly [42] – from simple question answering to highly structured agentic workflows – a significant gap remains in understanding agentic sophistication, viz., how to rank the many ways LLM capabilities can be employed in constructing agents, and the way such sophistication shapes LLM-based agents to replicate the variety of human strategic reasoning.
Not only is agentic sophistication essential but so too is the challenge of integrating diverse agentic AI systems within unified simulation environments. While several frameworks evaluate LLM-based agents through game-based interactions [47, 23] and specialised game-theoretic benchmarks [10, 27, 24, 17], no formally defined and widely adopted multi-agent simulation framework yet exists that can host heterogeneous agent architectures within a standardised environment. Instead, a variety of divergent benchmarking approaches have emerged, making it difficult to objectively determine which agents are most suitable for particular applications [19].
In addition to integration challenges, the black-box nature of LLMs introduces further concerns for agentic applications, particularly regarding reproducibility, explainability, and validation of agent behaviour [22]. These issues are especially acute in agent-based models designed for human-centered domains, where trustworthiness and safety are critical [21, 3]. Although several approaches have been proposed to address validation – such as formal representations, out-of-distribution analysis, and human-in-the-loop evaluation [7, 26, 25, 34] – many LLM-driven applications continue to lack systematic and rigorous validation mechanisms.
While these challenges remain underexplored and unresolved in game-theoretic social simulations, our work directly addresses them through the following contributions. We introduce guessing games as a testbed for strategic reasoning and present a human dataset, segmented by sub-populations, as a benchmark. Building on a formally defined multi-agent framework grounded in traditional agent concepts, we implement diverse agentic models and use this system to run agent simulations of two-player guessing games. Through these experiments, we explore the relationship between human-inspired sophistication and artificial reasoners, targeting two core problems: (1) replicating general patterns of human reasoning and (2) approximating individual-level variation. Our analysis spans over 2000 reasoning samples from 25 agent configurations, enabling detailed comparison of human and agent performance across multiple dimensions. Additionally, we introduce an obfuscated game scenario and an out-of-sample validation treatment to address potential biases from LLMs’ exposure to game-theoretic materials during training. Finally, we summarise our key findings and outline promising directions for future research.
2 Human Reasoning in 2-Player Guessing Games
Game theory provides a formal foundation for modelling strategic interactions among rational agents [30]. Guessing games, or p-Beauty Contests [20], are a canonical paradigm for evaluating players’ strategic reasoning, particularly their ability to anticipate others’ decisions [28]. In these games, $n$ players assumed strangers, simultaneously select a number from a defined range. The winner is the player whose choice is closest to $p$ times the mean of all selections.
A solution in guessing games is typically modelled using k-level reasoning theory [6], which assumes that each level-$k$ player believes others reason at most at level $k-1$ and best responds accordingly. This recursive process is formalised as: $a_{k}=p^{k}\cdot a_{0}$, where $a_{0}$ denotes the choice of 0-level players and $a_{k}$ the response of level-$k$ players. While in $n$-player repeated games players often iteratively reject weakly dominant strategies, gradually driving the average guess toward zero, the dynamics change considerably in two-player settings. Specifically, when $p<1$, choosing 00 becomes a strictly dominant strategy, as the lower guess will always win. Despite this theoretical simplicity, empirical studies show that players frequently select dominated strategies [28].
We focus on two-player one-shot guessing games, which offer several advantages. First, they remove the complexity associated with iterative learning and reasoning about future actions, thereby simplifying the analysis. Second, they likely reduce biases in large language models introduced by exposure to $n$-player beauty contest scenarios during pretraining. We adopt the two-player one-shot guessing game dataset from [28], which includes responses from two distinct participant cohorts:
  • Students: First-year undergraduates, assumed to have no prior exposure to game theory or strategic reasoning.
  • Experts: Professionals recruited at economics and psychology decision-making conferences, assumed to have relevant knowledge of game-theoretic concepts.
Figure 1: Distribution of human guesses per student and expert cohort. Vertical grey lines mark the standard k-levels for guessing games.
Theoretical expectations suggest that the expert group is more likely to identify the dominant strategy than students. This is reflected in our dataset, as $11.9\%$ of students and $37.7\%$ of experts chose 00. Furthermore, Figure 1 supports the expected intergroup performance difference: the expert cohort displays a substantially lower mean guess and greater density over lower values. Experts also exhibit higher skewness ($1.50$ vs. $0.55$). A Levene’s test found no significant difference in variances between the groups ($p=0.4351$), justifying the use of a standard independent samples $t$-test, which confirms a significant difference at $t=4.04$, $p=0.0001$.
3 Methods
3.1 Multi-agent Simulation Framework
A centralised multi-agent framework for agentic reasoning was introduced [38], combining an Observe–Orient–Decide–Act (OODA) control loop [39] with multi-level hypergames [43] to support recursive reasoning among agents. Here we define a streamlined variant that omits explicit evaluation of recursive reasoning and focuses on core agent interactions within our domain. We let $X$ denote the space of natural language game descriptions and $C$ the set of agent contexts. An umpire $\upsilon$, conceptualised as a pseudo-player [33], initialises each agent with a context $c_{i}\in C$ and instruction model $m_{i}\in M$. The umpire further enables gameplay by pairing agents and managing their participation in the guessing game, denoted as $G^{*}$.
Players’ internal processes are defined as the interpretation function $I:X\times C\times M\rightarrow\mathbb{Z}$ – wrapping the agent cycle’s orient and decision stages –, which we refer to as a cognitive process, producing a guess from a natural language game description, player context and instruction. Our experiments implement agents with different implementations of $I$.
The work in [38] entails a perspective-based formal definition of guessing games that captures individuals’ levels of recursive thinking. Our 2-player guessing games in this context are defined as $G=(N,p,A,U)$, where $N=\{i,j\}$ are our two players, and $p=\frac{2}{3}$ is the target multiplier. The action space is $A=A_{i}=A_{-i}\subseteq\mathbb{Z}$, where $i$ and $-i$ are the player and their opponent, $A$ is a finite interval of integers shared by both players, such as $A=\{0,1,\dotsc,100\}$ and $U:A\rightarrow\mathbb{R}$ is the utility function, specified as Equation 1 below.
$U_{i}(a_{i},a_{-i})=\begin{cases}1&\text{if }|a_{i}-p\mu|<|a_{-i}-p\mu|\\ 0.5&\text{if }|a_{i}-p\mu|=|a_{-i}-p\mu|\\ 0&\text{otherwise}\end{cases}$ (1)
where $\mu=\frac{a_{i}+a_{-i}}{2}$.
In the following sections, we introduce a set of agents within the framework defined above. First, we consider a game-theoretic model as a simple benchmark agent, then present two LLM-driven artificial reasoners that implement $I$ with different levels of conceptual sophistication.
3.2 Experience Weighted Attraction
Camerer et al. introduced the self-tuning experience weighted attraction model (EWA) as a single free-parameter game-theoretic learning algorithm designed to produce results that mimic human behaviour [16]. We implement the self-tuning EWA as a benchmark, given that it is a well-established state-of-the-art model from the economics literature. The heart of the model is the attraction mechanism composed as follows.
$\text{Memory}^{j}_{i}(t)=\phi\cdot N(t-1)\cdot\mathcal{A}^{j}_{i}(t-1)$ (2)
The memory component defines how past experience is carried forward, composed of the experience weight $N$ [5] and the change detector function $\phi$, defined in equations 3 and 4 below.
$N(t)=(1-\kappa)\cdot\phi\cdot N(t-1)+1$ (3)
$\phi(t)=1-\frac{1}{2}S_{i}(t)$ (4)
The surprise index function is composed of the immediate change vector $h$ and cumulative change vector $r$:
$S_{i}(t)=\sum^{m_{-i}}_{k=1}{(h^{k}_{i}(t)-r^{k}_{i}(t))^{2}}$ (5)
$h^{k}_{i}(t)=\frac{\sum^{t}_{\tau=1}{\mathcal{I}(s^{k}_{-i},s_{-i}(\tau))}}{t}$ (6)
$r^{k}_{i}(t)=\mathcal{I}(s^{k}_{-i},s_{-i}(t))$ (7)
where $\mathcal{I}$ is a binary index function.
The reinforcement component describes how new experiences are recorded:
$\text{Reinforcement}^{j}_{i}(t)=[\delta+(1-\delta)\cdot\mathcal{I}(s^{j}_{i},s_{i}(t))]\cdot\pi_{i}(s^{j}_{i},s_{-i}(t))$ (8)
The foregone payoffs are denoted by $\delta$:
$\delta_{ij}(t)=\begin{cases}1&\text{if }\pi_{i}(s^{j}_{i},s_{-i}(t))\geq\pi_{i}(t),\\ 0&\text{otherwise.}\end{cases}$ (9)
where $\pi_{i}(s^{j}_{i},s_{-i}(t))$ is $i^{\prime}s$ payoff for selecting strategy $j$ against $s_{-i}$ at time $t$. Players’ attractions are updated according to equation 10 below
$\mathcal{A}^{j}_{i}(t)=\frac{\text{Memory}^{j}_{i}(t)+\text{Reinforcement}^{j}_{i}(t)}{N(t-1)\cdot\phi\cdot(1-\kappa)}$ (10)
where $\kappa=0$, and players choose a strategy according to the distribution $P^{j}_{i}(t+1)=\frac{e^{\lambda\cdot\mathcal{A}^{j}_{i}(t)}}{\sum^{m_{i}}_{k=1}{e^{\lambda\cdot\mathcal{A}^{k}_{i}(t)}}}$, where $\lambda$ is the single parameter that requires manual tuning and controls the sensitivity of the model. The initial attraction $\mathcal{A}^{j}_{i}(0)$ is implemented according to the cognitive hierarchy model and the Poisson distribution function $P(k)=\frac{e^{-\tau}\tau^{k}}{k!}$ [6], and we set $N(0)=1$. Finally, we adopt $\lambda=2.39$ and $\tau=1.5$, which are the reported optimal parameters, tested on n-player guessing games [5].
Given that the self-tuning EWA model does not process natural language, we integrate EWA with our framework by implementing $I_{\text{EWA}}:G\rightarrow\mathbb{Z}$, where $I_{\text{EWA}}(G)\sim P(t+1)$ denotes sampling from the induced probability distribution over the strategies. In addition, in the case of EWA players, the responsibilities of $\upsilon$ are extended by the translation function $T:X\rightarrow G$, and the subject of game requests is $G^{*}$ in place of the natural language description $X$.
3.3 Artificial Reasoners
In this section, we introduce 3 categories of agentic and conceptual sophistication of the implementation suitably adapted from [38]: (1) the LLM used for reasoning, (2) the model used for composing the LLM prompts and (3) the agentic workflow the LLM is wrapped in.
LLM
We tested two state-of-the-art LLMs with our artificial reasoners [2] representing different levels of complexity: (1) Claude Haiku 3.5 is a quick and compact model with strong reasoning capabilities and (2) Sonnet 3.7 is the latest, most capable model, suitable for complex problem-solving.
Prompting Model
We construct our prompts based on the typical separation: context ($C$) and instruction ($M$). Providing LLMs with a character description aids their reasoning and instruction following [31, 24, 38]. We define three context configurations that represent different, increasingly more elaborate details for the LLM-driven agent as $C=\{c_{\varnothing},c_{\text{sim}},c_{\text{bio}}\}$:
  • No profile ($\mathbf{c_{\varnothing}}$): agents are presented with situational descriptions and are instructed to pick a number without further context;
  • Simple profile ($\mathbf{c_{\text{sim}}}$): agents are instructed to reason according to a role specification, whether the agent is an expert of game theory or a novice without any expertise;
  • Biography ($\mathbf{c_{\text{bio}}}$): the Umpire initialises each player with an elaborate background supporting the specific role.
Furthermore, we implement an additional configuration parameter based on a psychological model of human decision-making inspired by the agentic model of appropriateness (MoA) from [41]. We instruct our agents to address three questions in their responses: (1) What kind of situation is this?, (2) What kind of person am I? and (3)What should a person like me do in a situation like this? While the decoupled decision-making capability specified in our multi-agent framework already implements subtasking, having the LLM respond via the three questions integrates an additional layer of chain-of-thought prompting aiding reasoning [44]. Furthermore, as the three questions are associated with how humans make decisions, and as LLM training data contains a vast amount of samples of human culture, we adopt the hypothesis that guiding the agent through the human-inspired patterns will support the generated reasoning to better approximate human responses. We define MoA as a binary feature encapsulated in the instruction $M=\{m_{0},m_{1}\}$, where $m_{1}$ denotes the configuration that implements MoA.
Agentic Concept
First, implementing the decoupled revision and decision mechanism from [38], we define the Reasoner ($R$) agent as follows. The interpretation function $I:X\times C\times M\rightarrow\mathbb{Z}$ is defined as $I(x,c_{i},m_{i})=\delta(\rho(x,c_{i},m_{i}),c_{i},m_{i})$, where
  • $\rho:X\times C\times M\rightarrow\Xi$ is the reasoning function: $\rho(x,c,m)=\xi_{i}$, where $\xi_{i}\in\Xi$ is $i$’s belief about the opponent, a natural language reasoning based on game description $x$, context $c_{i}$, and instruction model $m_{i}$ on what the opponent guess $\hat{a}_{-i}$ is expected to be;
  • $\delta:\Xi\times C\times M\rightarrow A$ is the decision function: $\delta(\xi_{i},c_{i},m_{i})$, selecting $i$’s guess $a_{i}^{*}$ based on the belief $\xi_{i}$, context $c_{i}$, and instruction model $m_{i}$.
In contrast, the Simple ($S$) agent implements a one-step reasoning process, reflecting the growing association in recent literature between LLMs and agents through a looser definition of agency. Simple agents invoke $I(x,c_{i},m_{i})=\delta^{\prime}(x,c_{i},m_{i})$, where the decision function $\delta^{\prime}:X\times C\times M\rightarrow A$ internally applies a (possibly identity) mapping $f:X\rightarrow X^{\prime}$, producing a belief-enriched description $x^{\prime}\in X^{\prime}=X\times\Xi$. The resulting player guess $a_{i}^{*}$ may thus rely either on the original game description or on an inferred belief $\xi_{-i}$ about the opponent’s guess.
4 Experiments and Discussion
The following sections provide a detailed analysis of the experimental results, structured around multiple levels of behavioural fidelity.We evaluate agents’ ability to replicate human-like behaviour at two levels of granularity: first, at the population level, assessing how closely agent outputs align with aggregated human behaviour; and second, at the subpopulation level, measuring how well agents distinguish and replicate the behavioural differences between distinct human groups (e.g., students vs. experts). Within each level, we analyze performance across complementary dimensions of behavioural alignment: (1) mean response behaviour via $k$-level approximation; (2) distributional similarity, using Wasserstein distance; and (3) frequency of zero guesses. Finally, we extend our evaluation to an out-of-sample setting by modifying key game parameters to assess how well agents generalise beyond likely training data priors. For each pairwise comparison where mentioned, we first test for variance homogeneity using Levene’s test. We apply either a standard independent samples t-test (equal variances) or Welch’s t-test (unequal variances). If the data is severely skewed (|skewness| > 2), we use the non-parametric Mann–Whitney U test instead.
4.1 Benchmark Model Performance
Figure 2: Distribution of EWA over joint human results and k-levels.
As an established game-theoretic model, we first compare the self-tuning EWA agent with the overall human guess, as shown in Figure 2. The EWA model outperforms human results, averaging around $k=3$. Descriptive statistics confirm this observation: human guesses have a mean of $29.05$ and a standard deviation of $28.28$, while EWA agent guesses average $11.19$ with a standard deviation of $2.50$. Levene’s test indicates heterogeneous variances, and a Welch’s $t$-test confirms a statistically significant difference between the groups ($t=10.80$, $p<0.0001$). The Wasserstein distance between the two distributions is substantial ($W=22.34$). The EWA agent produces no zero guesses, indicating that while the EWA model achieves highly strategic behaviour, it systematically diverges from human reasoning patterns, failing to replicate key empirical properties of human play.
4.2 Artifical Reasoners
In the following sections, we test each LLM-based agent configuration across Haiku and Sonnet models, agent types – $S$ or $R$ – and variations of $C$ and $M$. We refer to specific agent configurations such that, e.g. $S_{c_{\varnothing},m_{0}}$ denotes the Simple agent with no profile, and no MoA or $R_{c_{bio},m_{1}}$ denotes the Reasoner agent with the Bio context using MoA. Given that EWA cannot process contextual information, we consider it implements $c_{\varnothing},m_{0}$. The two main test cases are specified by testing the (1) overall human-likedness of agents and (2) the ability to approximate human guesses specific to roles.
4.2.1 Population-wise Performance
First, we consider the case where the players’ individual details – i.e. the personal profile defining whether the player is a student or an expert – are not relevant. Data with subgroup representation is pooled together to represent the overall population.
In order to estimate and discuss the reasoning sophistication of our different implementations, we convert the guess values to k-levels as follows. Assuming irrational $k=0$ actors act randomly and produce a mean of 50 [6], we solve $(\frac{2}{3})^{k}\cdot 50=a_{i}^{*}$ to obtain $i$’s reasoning level. We set one dimension as agentic sophistication, where the order of sophistication is set as $\text{EWA}<S<R$. The second dimension describes the sophistication of the implemented $C$ and $M$, such that
$c_{\varnothing},m_{0}<c_{\varnothing},m_{1}<c_{\text{sim}},m_{0}<c_{\text{sim}},m_{1}<c_{\text{bio}},m_{0}<c_{\text{bio}},m_{1}$
defines the order of sophistication. Our initial expectation is that there is a linear relation between the two axes of sophistication and the human-likedness of the results. In Figure 3, we present LLM-wise heatmaps of the absolute mean error (AME) $\left|\bar{k}_{\text{true}}-\bar{k}_{\text{pred}}\right|$ of the true human and the agent predicted k values. We can observe that the least complex implementation (EWA) showed the most distant performance, and Haiku produced the best results with the most complex implementation, which matches humans’ reasoning level with a difference of just $0.02$. However, while the opposite ends of the complexity range matched our expectations, the intermediate results are not always consistent with the progressive inclusion of more sophisticated implementations. Sonnet shows the opposite reaction to more advanced configurations and, overall, is more distant from human reasoners than the simpler LLM.
We transform our categorical features into ordinal variables as follows: agent $\in\{\text{EWA}=0,S=1,R=2\}$, $c\in\{c_{\varnothing}=0,c_{sim}=1,c_{bio}=2\}$, $m\in\{m_{0}=0,m_{1}=1\}$, and model $\in\{\text{Haiku}=0,\text{Sonnet}=1\}$. Then, the overall sophistication score is computed as their sum. Table 1 reveals the statistical effects of individual sophistication factors on our LLM-based agents, measured by linear regression models on the k-level AME. While the agent type does not significantly influence the results, the sophistication of LLM, $C$ and $M$ have significant effects. However, the effect direction is not aligned with our expectations, and a Spearman correlation test confirms a weak and insignificant negative correlation between the sophistication score and deviation from humans’ k-level.
Figure 3: Heatmap showing the absolute distance between agents’ and humans’ mean reasoning sophistication according to k-levels.
While the corresponding k-levels are convenient and meaningful measures of the overall performance, they are not informative enough for individual-based simulations where the goal is to match the distribution of human guesses. We calculate the kernel density estimates and use the Wasserstein distance to measure the similarity between the distributions of human and agent guesses. The distributions are presented in Figure 4. Table 2 summarises the Wasserstein distances between each agent configuration per LLM and human guesses and Table 1 contains the statistical effects of sophistication factors on the distance. The distribution-related results paint a slightly different picture than the population mean-derived metrics. While, overall, Haiku still outperforms Sonnet in matching human results, and the best match is achieved by $R$, the most human-like result is produced by the least sophisticated $c_{\varnothing},m_{o}$ configuration. Similarly to before, while Haiku benefits from the more sophisticated agent implementation, Sonnet performs better with the simpler alternative. Another key difference between the two models is that the inclusion of MoA had a positive effect on Haiku in most configurations with both $S$ and $R$, but it had little to no effect on Sonnet’s reasoning.
Figure 4: KDEs of $S$ (top) and $R$ (bottom) agents with $C$,$M$ variations. Red and blue correspond to overall results, light/dark shades correspond to cohorts.
Table 1: Statistical effects of sophistication factors and overall sophistication score on $k$-level error, Wasserstein distance, and deviation from human zero-guessing rate.
Factor $k$-Level Error Wass. Distance 0-Rate Error
Coef. p-val Coef. p-val Coef. p-val
Agent Type 0.073 0.8822 -0.452 0.6484 -2.609 0.1734
LLM Model 2.524 0.0000 0.416 0.6748 -2.609 0.1734
$C$0.564 0.0186 -1.543 0.2012 -2.586 0.2852
$M$ 1.272 $0$1.033 0.2926 -1.442 0.4585
Spearman $\rho$0.025 0.1901 -0.055 0.7987 -0.042 0.8447
Table 2: Wasserstein distances between agents and humans for each mind type across LLMs. Bold values indicate the overall best match per mind (smallest distance), italics mark the best match per LLM, and asterisks highlight the best score in the table.
Config EWA Simple Reasoner
Haiku Sonnet Haiku Sonnet
$c_{\varnothing},m_{0}$ 22.34 22.05 18.59 14.28* 20.38
$c_{\varnothing},m_{1}$ N.A. 16.69 19.79 17.51 20.88
$c_{\text{sim}},m_{0}$ N.A. 19.92 15.41 17.46 19.32
$c_{\text{sim}},m_{1}$ N.A. 15.05 14.38 17.42 18.88
$c_{\text{bio}},m_{0}$ N.A. 22.99 16.57 18.37 20.33
$c_{\text{bio}},m_{1}$ N.A. 19.08 16.24 16.14 20.78
The human data indicates an overall rate of $23.15\%$ of players managed to pick the equilibrium strategy: 00. While EWA produced the highest mean k-level, indicating a bias towards optimisation instead of human-level performance, it still fails to derive the dominant strategy. In the case of the LLM-based agents, our general observation so far was that game-theoretic problem solving is more dominant in Sonnet’s results than in Haiku’s. While by the previous metrics, this had a negative effect on Sonnet’s ability to replicate human results, the current results show it helps with reaching zero guesses – shown in Table 3 – and zero choices appear in most settings. Haiku is unable to replicate zero guesses in most cases, with $S_{c_{\text{sim},m_{1}}}$ as the sole exception at $14\%$. While MoA improves the rate of 00 guesses for $S$ with Sonnet and both models’ best results are associated with MoA, the overall sophistication score does not significantly correlate with the agents’ performance of matching humans’ 00 guessing rate.
Table 3: 00 guesses per 100 agents (Sonnet only). Human benchmark at $23$.
Agent $\mathbf{c_{\varnothing},m_{0}}$ $\mathbf{c_{\varnothing},m_{1}}$ $\mathbf{c_{\text{sim}},m_{0}}$ $\mathbf{c_{\text{sim}},m_{1}}$ $\mathbf{c_{\text{bio}},m_{0}}$ $\mathbf{c_{\text{bio}},m_{1}}$
$S$ $2$ $52$ $5$ $10$ $9$ $11$
$R$ $4$ $8$ 00 00 $1$ $1$
4.2.2 Subpopulation-wise Performance
Simulations that require an individual-based approach where specific subpopulations can be represented explicitly must implement role specification. The following analysis concerns the configurations supporting profile assignment and investigates how well artificial agents can replicate the intergroup differences, role-specific reasoning levels, and guess distributions. Figure 1 depicting human results per cohort shows a higher density of guesses over lower values and a significantly lower mean of guesses for experts. In this section, we use those results as a benchmark and evaluate artificial agents’ ability to replicate the significant difference in reasoning between human students and experts.
First, we calculate the k-level means for students and experts separately and define $\Delta=\left|\bar{k}_{\text{true}}^{\text{student}}-\bar{k}_{\text{pred}}^{\text{student}}\right|-\left|\bar{k}_{\text{true}}^{\text{expert}}-\bar{k}_{\text{pred}}^{\text{expert}}\right|$ to measure if the agent replicates student and expert behaviour equally well/poorly, where values close to 00 indicate an approximation of the difference between the human groups. Figure 5 entails the $\Delta$ heatmaps for Haiku and Sonnet. The dominantly negative values in the case of $S$ agents across both models indicate that these agents are consistently closer to expert behaviour, where likely the overestimation of student reasoning levels is due to an overfit to game-theoretic optimizations we observed in previous instances. On the other hand, $R$ agents show a more balanced behaviour and manage to approximate better the human differences with all $C\times M$ errors within +/0.4+/-0.4+ / - 0.4 k levels. In addition, the transition from $m_{0}$ to $m_{1}$ consistently causes agents to optimise more and overestimate reasoning levels. The statistical significance test indicates lesser divergence in the case of $R$ agents using Haiku and the minor and insignificant differences in the case of $R_{c_{\text{sim}},m_{0}}$ and $R_{c_{\text{bio}},m_{0}}$ indicate closer matching of the human intergroup divergence.
Figure 5: Delta between agents’ and humans’ cohort-wise mean k-level difference with significance levels: * p < 0.05, ** p < 0.01, *** p < 0.001.
Figure 4 is extended with the profile-specific distribution curve per model for the configurations where the distinction between cohorts is available, and Table 4 collects the Wasserstein distances for each agent variation.
Table 4: Wasserstein distances between agent and human guess distributions for each agent type across LLMs. Bold values indicate the overall best match per agent (smallest distance) for the given cohort; italics mark the best match per LLM, and * highlights the best score in the table.
Config Student Expert
Simple Reasoner Simple Reasoner
H S H S H S H S
$c_{\text{sim}},m_{0}$ 16.82 18.6 18.47 20.70 23.65 16.37 18.60 17.40
$c_{\text{sim}},m_{1}$ 18.74 16.98 18.16 21.04 14.68 15.84 17.33 18.02
$c_{\text{bio}},m_{0}$ 20.39 18.63 18.01 21.50 25.36 14.31* 18.82 18.28
$c_{\text{bio}},m_{1}$ 16.10 16.93 15.69* 23.14 22.69 15.59 17.22 17.54
In the case of Haiku, we observe that the inclusion of MoA has an overall negative correlation with the Wasserstein distance. Similar to the previous set of evaluations, the effects of the different configurations and the more complex reasoning model on Sonnet are inconsistent; no well-defined patterns emerge. The relative overoptimization of Sonnet and its poorer instruction following ability leads to a similar resistance to increasing complexity as we have seen in the experiments without profile specification. In 3/4 cases, Haiku and the $c_{\text{sim},m_{1}}$ configuration achieved the closest results. In the case where Haiku did worse, it produced an outlier distribution relative to its overall performance, which has a significantly lower variance and higher mean. The statistics show a high number of agents choosing 22 and 33, which are perfectly aligned with the $k=1$ and $k=2$ levels associated with guessing games. A manual random-sampled analysis of the produced reasoning samples indicates that the specific configuration seemingly caused the agent to pick up on the corresponding game-theoretic material. An in-depth analysis of the corresponding natural language reasoning samples may reveal further insights and is needed to validate our suspicions, but lies outside the scope of this paper.
We investigate how the zero guesses are spread over the subpopulations. The sparsity of zero guesses remains, but a consistent pattern we can observe is that, according to our expectation, expert agents guess zero more frequently than student agents, which mirrors human behaviour, where experts were thrice as likely to guess zero.
[…]
Factor Student Expert
$k$-Err. Was. Dist. 0 Err. $k$-Err. Was. Dist. 0 Err.
Agent $-3.80^{***}$ $-1.69$ $0.25$ $-6.90^{***}$ $0.66$ $-12.00^{**}$
LLM $0.97^{***}$ $1.89$ $-0.25$ $2.46^{***}$ $-3.13^{*}$ $-5.50$
$C$ $-0.10$ $-0.11$ $0.25$ $0.73^{*}$ $-0.99$ $-2.00$
$M$ $-0.03$ $-0.79$ $0.25$ $1.51^{***}$ $-1.73$ $-5.50$
$\rho$ $0.19^{***}$ $0.01$ $-0.18$ $-0.08^{*}$ $0.03$ $0.21$
The group-specific regression analysis in Table 5 supports our earlier observations. Despite the sparsity of zero-guessing behaviour, we find that Reasoner agents ($R$) align significantly better with expert zero-guess rates. For distributional similarity (Wasserstein distance), most factors have no significant effect; the exception is the LLM model choice, which exhibits a significant effect for experts and a near-significant trend for students ($p=0.07$), with opposite signs across the two groups. Agent type and LLM model have a strong and consistent influence on $k$-level approximation for both cohorts: more complex reasoning agents using the smaller LLM (Haiku) achieve lower error. In contrast, contextual ($C$) and instruction complexity ($M$) have only minor and non-significant effects for students, while for experts, increased complexity correlates with higher $k$-level error. Finally, the Spearman $\rho$ correlation analysis reveals a significant but modest association between overall sophistication and $k$-level error: positive for students, suggesting over-complexity harms alignment, and negative for experts, indicating modest improvements. No monotonic trends emerge for the other two metrics.
4.3 Out-of-Sample Validation
Most guessing game experiments involve $n$ players selecting numbers between 00 and $100$ [29]. In contrast, our 2-player variant may be underrepresented or entirely absent from LLM training. We treat our setup as an out-of-sample case and further examine generalization by shifting key parameters to diverge from training data priors.
Shifted Range
Humans are expected to understand the game regardless of varying parameters [8]. We test whether artificial agents can do the same by shifting the range of valid guesses to $[100,200]$, following [15]. This allows us to assess whether our artificial reasoners are misled by their training data where this problem is underrepresented or can identify the distinct game parameters and adapt their strategy to novel numerical boundaries.
As with the standard range, the smaller model, Haiku, outperforms Sonnet in its ability to generalise for the modified parameters and provide valid responses – as depicted by Figure 6. We can also observe a trend showing that simpler implementations can generalise better. In addition to the performance gap between Haiku and Sonnet, the transition from $S$ to $R$ also decreases the scores. The inclusion of MoA results in a drop in accuracy in each case. Except for the Simple agent using Haiku, the increasing complexity in $C$ shows a negative correlation with guess validity.
Figure 6: Trend of ratios of valid guesses across increasingly complex configs and agent models.
Assuming human understanding of the game is invariant to the number range, we normalise shifted-range guesses by subtracting 100 and compare them to standard-range distributions. We then compute Wasserstein distances and t-statistics between original and scaled conditions to identify robustness across agent variations.
Table 6: Wasserstein distances (W) and t-statistics (t) comparing KDEs for original vs. scaled guesses. Asterisks mark statistically significant t-values ($p<0.05$). The closest significant matches per agent type are in bold.
Config Simple Reasoner
Haiku Sonnet Haiku Sonnet
W t W t W t W t
$c_{\varnothing},m_{0}$ 2.70 0.02 15.99 12.34* 7.69 2.77* 13.20 18.36*
$c_{\varnothing},m_{1}$ 9.06 2.10* 9.26 6.39* 13.62 5.87* 27.84 -0.55
$c_{\text{sim}},m_{0}$ 7.21 4.70* 12.19 8.61* 7.60 1.54 13.63 14.41*
$c_{\text{sim}},m_{1}$ 8.81 -4.06* 12.64 8.33* 14.76 7.00* 14.78 5.12*
$c_{\text{bio}},m_{0}$ 5.48 5.46* 14.05 12.08* 8.39 1.04 12.29 15.03*
$c_{\text{bio}},m_{1}$ 7.34 4.38* 13.79 11.80* 13.38 6.31* 12.36 26.41*
The results in Table 6 reinforce earlier findings: Haiku does better than Sonnet, and $R$ agents are more dissimilar. While increasing the $C$ complexity does not lead to significant and consistent changes, $m_{1}$ increases the distance score in each case over $m_{0}$.
4.4 Cost of Complexity
Given that $R$ agents involve an additional round of interaction with the LLM per player and that the MoA and the increased amount of details in the base configurations cause lengthier requests and potentially longer responses, in this section, we summarise the estimated incurred costs in terms of input and output token counts in Table 7.
Table 7: Cost of LLM agent sophistication in terms of tokens.
$c_{\varnothing},m_{0}$ $c_{\varnothing},m_{1}$ $c_{\text{sim}},m_{0}$ $c_{\text{sim}},m_{1}$ $c_{\text{bio}},m_{0}$ $c_{\text{bio}},m_{1}$
S In 106.3 166.0 161.8 200.9 410.5 447.9
Out 189.3 311.3 256.1 414.5 243.0 330.1
R In 292.9 390.0 341.7 642.4 878.7 935.6
Out 503.7 632.9 530.8 666.1 547.9 649.0
5 Conclusions
We conducted an in-depth analysis of the relationship between agentic sophistication and human-likeness in strategic reasoning comparing LLM-driven artificial reasoners with human-data from 2-player guessing games. We used a novel multi-agent-based simulation framework with high conceptual integrity to host agents with varying levels of conceptual complexity representing both looser and stronger agentic definitions, integrating state-of-the-art compact and flagship LLMs. We demonstrated that human-inspired cognitive architectures can successfully mimic human strategic reasoning, but the relationship between implementation complexity and human-likeness is non-linear and is objective-dependent. In addition, our out-of-sample validation with the shifted range of guesses revealed that the compact LLM and simpler configurations generalised better. The evaluation of both main objectives and the robustness test provided consistent insights, which suggest that the larger language model is likely more exposed to overfitting and, thus, generalises worse. These results suggest that researchers should identify the correct models, concepts and architectures to match their specific use case requirements. Defaulting to convenient, simple agentic concepts can lead to suboptimal results; similarly, highly conceptualised implementations may sidetrack agents from the target behaviour.
Promising directions for future research that emerge from our work involve expanding the test beyond a guessing game and evaluating the generalizability of agents across different game-theoretic scenarios. In addition, an in-depth semantic analysis of agents’ reasoning processes and a comparison of human and agent natural language reasoning traces and emergent patterns will provide valuable insights into the observed effects of increased complexity and will help researchers design objective-specific artificial reasoning agents.
{ack}
The first author would like to express his gratitude to David Levine for his valuable insights.
This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001).